OpenStudy (rebecca1588):

Is each situation a linear or a nonlinear model? A home’s value is given by the equation y = 154,000(1.04)x, where y is the value of the house and x is the number of years since the home was purchased. A. linear model B. nonlinear model

OpenStudy (anonymous):

it is definitely not a line a line looks like \(y=mx+b\)

OpenStudy (anonymous):

what you have there is an "exponential" model the variable is in the exponent

OpenStudy (rebecca1588):

@Directrix

Directrix (directrix):

Is this >> y = 154,000(1.04)x supposed to be y = 154,000(1.04)^x ?

OpenStudy (rebecca1588):

Yeah sorry :)

Directrix (directrix):

It is not linear. If it were, the x would appear as x to the first power.

OpenStudy (rebecca1588):

Okay Thanks :) I only have two more now.

Directrix (directrix):

The graph of this: y = 154,000(1.04)^x is not a line.

OpenStudy (rebecca1588):

Ooh okay

Directrix (directrix):

Linear looks like y = 5x + 2. The big deal is that the x is not raised to a power such as 2 and the x is not an exponent.

Directrix (directrix):

You can always graph the function and look and see if the graph is a line.

OpenStudy (rebecca1588):

Oh ok I think my last two might be linear then @Directrix 1. The amount of money Elijah has in his bank account is given by the equation B = 1400 + 120m, where B is the balance in his account and m is the number of months since he opened the account. 2. The number of words Sarah can type in a given amount of time is given by the equation w = 32m, where w is the number of words and m is the number of minutes she types.

OpenStudy (rebecca1588):

1 looks like it, and maybe 2?

Directrix (directrix):

I agree with you that both are linear. No exponents, no powers.

OpenStudy (rebecca1588):

Yay thanks so much for teaching how do these :)